# Privacy Beyond Pixels: Latent Anonymization for Privacy-Preserving Video Understanding

### v1

@@ 1 Introduction
Video foundation models (VFMs) capable of addressing multiple video understanding tasks wang2022internvideo; wang2023videomaeV2; vjepa; dave2024nomore have been developed, but due to powerful modeling capabilities, the visual features extracted by these models reveal sensitive private information such as skin color, clothing, or gender (Figure 1). The high quality spatio-temporal features have enabled usage in real-world scenarios such as patient monitoring, sports analytics, robotics, and surveillance, where it is common practice to extract visual features, store them, and utilize them across multiple tasks. However, an attacker can use a classifier on these features to disclose such private attributes fioresi2023ted, making it unsafe to store or share the visual features directly. This raises the question of how to protect an individual’s private information in the feature space while maintaining the powerful video understanding capabilities of the VFMs.
Prior privacy-preserving methods wu2020privacy; dave2022spact; li2023stprivacy; fioresi2023ted anonymize video models at the expense of utility. These works have primarily focused on input-level (i.e., pixel-level) privacy, which has limited applicability for two main reasons: (1) Since it alters the input, it requires retraining the utility model on the transformed data, which is impractical for large pretrained models trained on millions of videos with specific training recipes. (2) Existing methods have only proven efficacy on single downstream tasks. For instance, SPAct dave2022spact is suitable only for action recognition, whereas TeD-SPAD fioresi2023ted is limited to anomaly detection. Consequently, the existing pixel-level anonymization formulation is not suitable to adopt to new developments in utility video foundation models.
@@
To anonymize the utility model latent space, we introduce a lightweight, learnable Anonymizing Adapter Module (AAM) on a frozen VFM. Importantly, in contrast to prior methods wu2020privacy; dave2022spact; fioresi2023ted, AAM is applied to temporal clip-level features instead of individual frames, allowing the anonymizer to communicate across the temporal dimension, more naturally aligning with video tasks. Our training framework employs three key objectives: (1) a clip-level self-supervised privacy preservation objective to minimize mutual information between two static clips, (2) a co-training utility objective to maintain performance across predefined tasks, and (3) a latent consistency loss to ensure generalization on unseen tasks. SPLAVU integrates seamlessly with multiple state-of-the-art methods for downstream tasks, significantly outperforming previous privacy-preserving methods. Additionally, SPLAVU is data-efficient; even when trained on a small dataset like HMDB51 kuehne2011hmdb, it generalizes effectively without compromising the privacy-utility tradeoff. Under the right conditions, it can even defend against recognition of motion-based sensitive attributes such as gait. Beyond privacy protection, our work addresses the emerging issue of human-attribute bias in video understanding. For instance, models may exhibit gender biases by associating certain actions with specific genders based on stereotypes. For the first time, we introduce protocols to evaluate and mitigate gender bias in action recognition models.
@@ 2 Related Work
Video understanding spans tasks like action recognition, temporal action localization, and weakly-supervised anomaly detection. Various datasets have been introduced  carreira2017quo; hvu; goyal2017something; hacs, and recent advancements include self-supervised jenni2021time; dave2024nomore; dave2022tclr; thoker2023tubelet and foundational models bardes2024revisiting; wang2023videomaeV2; wang2022internvideo capable of handling multiple video understanding tasks, enhancing versatility and generalizability.
@@
Privacy Preservation in Video Understanding Recent efforts in video action recognition have addressed visual privacy concerns. Many studies have aimed to protect visual privacy at the time of data capture by utilizing non-intrusive sensors such as thermal imaging, depth cameras, and event cameras luo2018computer; hinojosa2022privhar; kim2022privacy; ahmad2022event; ahmad2023person. In this study, we focus exclusively on models using standard RGB cameras. Initial approaches involved reducing the resolution of input data ryoo2017privacy; dai2015towards; liu2020indoor or employing object detection for targeted obfuscations ren2018learning; zhang2021multi. However, recent research indicates that these methods often fail to balance utility and privacy effectively wu2020privacy; dave2022spact; fioresi2023ted; kumawat2022privacy; peng2024joint; li2023stprivacy. Wu et alwu2020privacy showcased an adversarial training framework where a U-Net ronneberger2015u modifies input frames to decrease the accuracy of private attribute prediction while preserving action recognition performance.  dave2022spact proposed a self-supervised variant that focuses on reducing mutual information instead of relying on sensitive privacy labels.  fioresi2023ted adapts the self-supervised privacy objective from dave2022spact for the anomaly detection task. Compared to the prior input-level anonymization methods our latent-space anonymization method differs in two key aspects: (1) previous methods are tailored to specific downstream tasks, such as action recognition in dave2022spact and anomaly detection in fioresi2023ted, while our approach aims to preserve privacy across various downstream video understanding tasks, (2) unlike prior methods, our method does not necessitate the retraining of the video model, thus providing computational efficiency for anonymizing even large-scale video foundation models.
@@
Bias Mitigation Computer vision tasks often struggle with spurious correlations geirhos2018imagenet; geirhos2020shortcut, where models rely on irrelevant information to make decisions, such as using background cues for action recognition instead of focusing on subjects’ movements ding2022motion; li2023mitigating; fioresi2025albar. Unfortunately, biases across a variety of protected demographic attributes, such as perceived gender, skin color, and age zhao2017men; stock2017convnets; buolamwini2018gender; wilson2019predictive; tong2020investigating; steed2021image; gustafson2023facet; narnaware2025sb have been found in vision-based tasks. These biases not only skew model performance but can also perpetuate harmful stereotypes. barbano2021bridging explored the relationship between debiasing and privacy preservation, finding that there exists a subset of privacy preservation methods that are suitable for debiasing, giving promise to privacy preservation as a form of debiasing. In contrast to the image domain, biases in the video domain have not been as extensively studied. While a few papers choi2019can; li2023mitigating address and mitigate scene bias in action recognition tasks, they overlook biases related to human attributes. Motivated by this gap, we introduce, for the first time, protocols to assess gender bias in action recognition. Our findings demonstrate that our self-supervised privacy preservation method, even without an explicit bias-related objective, effectively generalizes in mitigating gender bias.
@@ 3.1 Problem Formulation
Privacy Preservation We propose a novel privacy-preserving framework that handles multiple utility tasks across diverse video datasets. Our framework is designed to maintain the high performance of a frozen video encoder across tasks while enforcing robust privacy constraints. Specifically, we consider video datasets that span action recognition ($\mathbb{D}_{reco}$), temporal action detection ($\mathbb{D}_{tad}$), and anomaly detection ($\mathbb{D}_{anomaly}$). Each dataset $\mathbb{D}$ contains $N$ video samples, represented as $\{{\mathbf{x}^{(i)},\mathbf{y}^{(i)}\}_{i=1}^{N}}$, where $\mathbf{x}^{(i)}$ is a video instance, and $\mathbf{y}^{(i)}$ is its corresponding task-specific label. We define the set of utility tasks as $\{T_{AR};T_{TAD};T_{AD}\}\in T^{*}$. In implementation, we can choose a subset of tasks from $T^{*}$ for training, then evaluate on held-out tasks. We introduce a budget privacy evaluation task, denoted as $B$, where performance is measured by private attribute prediction. The framework starts with an off-the-shelf video encoder model $f_{E}$, left completely frozen. The overall goal of $f_{A}$ is threefold: (1) to maintain the performance of $f_{E}$ across the set of defined utility tasks $T^{*}$, (2) to simultaneously reduce the performance on budget private attribute prediction task $B$, and (3) to preserve the general capabilities of $f_{E}$ such that performance is maintained on unseen tasks. This privacy preservation framework is outlined via the following criteria:
Criterion-1: Across each utility task, performance should be retained. Specifically, for task $T^{n}$, the loss $\mathcal{L}_{T^{n}}$ before and after anonymization should be approximately equal.
$\begin{array}[]{l}\sum^{|T^{*}|}_{n}(\mathcal{L}_{T^{n}}(f_{T^{n}}(f_{A}(f_{E}(X))),Y),T^{n})\\ \approx\sum^{|T^{*}|}_{n}(\mathcal{L}_{T^{n}}(f_{T^{n}}(f_{E}(X)),Y),T^{n}).\end{array}$ (1)
@@
Perceived Gender Bias Evaluation In the standard bias evaluation protocol, we are given a video dataset $\mathbb{D}_{reco}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}^{N_{IID}}_{i=1}$, where $\mathbf{x}^{(i)}$ is the $i$th video instance, $\mathbf{y}^{(i)}$ is the associated action label, and $N_{IID}$ is the number of dataset instances. After training, performance is evaluated on unseen bias test set $\mathbb{D}_{reco-OOD}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}^{N_{OOD}}_{i=1}$, where $N_{OOD}$ is the number of out-of-distribution instances. The aim of any debiasing technique is to learn generalizable features of $\mathbb{D}_{IID}$ such that performance is maximized on $\mathbb{D}_{OOD}$ without compromising IID performance.
When considering perceived gender information, our in-distribution video dataset is now formulated as $\mathbb{D}_{reco}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)},\mathbf{g}^{(i)})\}^{N_{IID}}_{i=1}$, where $\mathbf{g}^{(i)}\in\{male,female\}$ is the perceived binary gender label. We acknowledge that this binary formulation is not ideal and not inclusive of all gender categories. The bias evaluation test set also includes label $\mathbf{g}$ in order to evaluate subclass performance. Final evaluation is on a test set with a different label distribution.
@@ 3.2 Anonymization Framework
This section describes the full anonymization framework and training. The framework consists of 3 major components: (1) a frozen video encoder backbone $f_{E}$, (2) an anonymization function adapter $f_{A}$, which modifies the latent features while retaining the original shape, and (3) a set of utility classifier heads $\{f_{T_{AR}};f_{T_{TAD}};f_{T_{AD}}\}\in f_{T^{*}}$ for a predefined set of tasks.
Network Initialization To start, we initialize $f_{A}$ to act as an identity function. The video encoder model $f_{E}$ is initialized with off-the-shelf weights of Kinetics400 carreira2017quo pretraining. Each $f_{T}$ classifier head matches a standard architecture for the provided task. For stability, these are initialized through non-anonymized training on their respective utility tasks. For action recognition, $f_{T_{AR}}$ is a simple linear layer. For action detection $f_{T_{TAD}}$ and anomaly detection $f_{T_{AD}}$ architectures, TriDet shi2023tridet and MGFN chen2023mgfn respectively are utilized.
Anonymization Training
The training process consists of an adversarial optimization between a budget privacy loss $\mathcal{L}_{B}$ and a collection of standard utility losses $\mathcal{L}_{T^{*}}$, regularized by a proposed latent consistency loss $\mathcal{L}_{LC}$.
@@
Clip-Level Budget Privacy Objective Our clip-level self-supervised budget privacy objective is the key component for facilitating anonymization without requiring private attribute labels. The intuition is that two frames share a lot of mutual information, so if we minimize the similarity between them, the shared spatial information gets destroyed. A crucial difference setting SPLAVU apart from prior works is that the anonymizer works across the temporal dimension using 3D clip features instead of a 2D U-Net wu2020privacy; dave2022spact; fioresi2023ted. This way, when combined with utility task losses, the anonymization model learns to remove all spatial information, maintaining only temporal information useful for solving the utility task. We utilize the SimCLR NT-Xent simclr contrastive loss as $\mathcal{L}_{B}$, defined as follows:
@@
where $\mathbf{\bar{h}}_{t}^{(i)}$ represents the feature vector of a static clip sampled from video $\mathbf{x}^{(i)}$ at time $t$, $d(u,v)=exp(u^{T}v/(\|u\|\|v\|\tau))$ computes the similarity between the input vectors with temperature parameter $\tau$. $\mathds{1}_{[j\neq i]}$ is an indicator function that equals 1 when $j\neq i$. Minimizing this loss increases the similarity between inputs $\mathbf{\bar{h}}_{\bar{t_{1}}}^{(i)}$ and $\mathbf{\bar{h}}_{\bar{t_{2}}}^{(i)}$ over the sum of all other clips in the batch. Instead, we opt to maximize the loss, resulting in the objective destroying mutual information between these clips instead. Notably, as opposed to previous works wu2020privacy; dave2022spact; fioresi2023ted, we do not use a disjoint image encoder model to compute $\mathbf{\bar{h}}_{\bar{t}}^{(i)}$. Alternatively, the video encoder model $f_{E}$ itself is used to process clip frames (tiled to match a standard clip shape, see Figure 2), leading to a much more natural utility-privacy interaction for improved anonymization. These static clip features are then directly utilized in this budget privacy loss (Eq. 5).
@@
where $\omega_{R}$, $\omega_{T}$, and $\omega_{B}$ are weights to control the strength of each objective. The privacy loss $\mathcal{L}_{B}$ works against $\mathcal{L}_{T}$ and $\mathcal{L}_{LC}$ in a GAN-style paradigm, until the anonymizer is able to remove all encoded spatial information except for what is necessary for performance on the utility tasks. After this training, we are left with a lightweight anonymization adapter $f_{A}$ that can be appended to the off-the-shelf video encoder model $f_{E}$ for use in a variety of downstream tasks.

### v2

@@ 1 Introduction
Video foundation models (VFMs) capable of addressing multiple video understanding tasks Wang et al. (2022; 2023); Bardes et al. (2024b); Dave et al. (2024) have been developed, but due to powerful modeling capabilities, the visual features extracted by these models reveal sensitive private information such as skin color, clothing, or gender (Figure 1). The high quality spatio-temporal features have enabled usage in real-world scenarios such as patient monitoring, sports analytics, robotics, and surveillance, where it is common practice to extract visual features, store them, and utilize them across multiple tasks. However, an attacker can use a classifier on these features to disclose such private attributes Fioresi et al. (2023), making it unsafe to store or share the visual features directly. This raises the question of how to protect an individual’s private information in the feature space while maintaining the powerful video understanding capabilities of the VFMs.
Prior privacy-preserving methods Wu et al. (2020); Dave et al. (2022b); Li et al. (2023b); Fioresi et al. (2023) anonymize video models at the expense of utility. These works have primarily focused on input-level (i.e., pixel-level) privacy, which has limited applicability for two main reasons: (1) Since it alters the input, it requires retraining the utility model on the transformed data, which is impractical for large pretrained models trained on millions of videos with specific training recipes. (2) Existing methods have only proven efficacy on single downstream tasks. For instance, SPAct Dave et al. (2022b) is suitable only for action recognition, whereas TeD-SPAD Fioresi et al. (2023) is limited to anomaly detection. Consequently, the existing pixel-level anonymization formulation is not suitable to adopt to new developments in utility video foundation models.
@@
To anonymize the utility model latent space, we introduce a lightweight, learnable Anonymizing Adapter Module (AAM) on a frozen VFM. Importantly, in contrast to prior methods Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023), AAM is applied to temporal clip-level features instead of individual frames, allowing the anonymizer to communicate across the temporal dimension, more naturally aligning with video tasks. Our training framework employs three key objectives: (1) a clip-level self-supervised privacy preservation objective to minimize mutual information between two static clips, (2) a co-training utility objective to maintain performance across predefined tasks, and (3) a latent consistency loss to ensure generalization on unseen tasks. SPLAVU integrates seamlessly with multiple state-of-the-art methods for downstream tasks, significantly outperforming previous privacy-preserving methods. Additionally, SPLAVU is data-efficient; even when trained on a small dataset like HMDB51 Kuehne et al. (2011), it generalizes effectively without compromising the privacy-utility tradeoff. Under the right conditions, it can even defend against recognition of motion-based sensitive attributes such as gait. Beyond privacy protection, our work addresses the emerging issue of human-attribute bias in video understanding. For instance, models may exhibit gender biases by associating certain actions with specific genders based on stereotypes. For the first time, we introduce protocols to evaluate and mitigate gender bias in action recognition models.
@@ 2 Related Work
Video understanding spans tasks like action recognition, temporal action localization, and weakly-supervised anomaly detection. Various datasets have been introduced  Carreira and Zisserman (2017); Diba et al. (2020); Goyal et al. (2017b); Zhao et al. (2019), and recent advancements include self-supervised Jenni and Jin (2021); Dave et al. (2024; 2022a); Thoker et al. (2023); Swetha et al. (2021) and foundational models Bardes et al. (2024a); Wang et al. (2023; 2022); Swetha et al. (2023) capable of handling multiple video understanding tasks, enhancing versatility and generalizability.
@@
Privacy Preservation in Video Understanding Recent efforts in video action recognition have addressed visual privacy concerns. Many studies protect visual privacy at the time of data capture by utilizing non-intrusive sensors such as thermal imaging, depth cameras, and event cameras Luo et al. (2018); Hinojosa et al. (2022); Kim et al. (2022); Ahmad et al. (2022; 2023). In this study, we focus exclusively on models using standard RGB cameras. Initial approaches involved reducing the resolution of input data Ryoo et al. (2017); Dai et al. (2015); Liu and Zhang (2020) or employing object detection for targeted obfuscations Ren et al. (2018); Zhang et al. (2021). However, recent research indicates that these methods often fail to balance utility and privacy effectively Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023); Kumawat and Nagahara (2022); Peng et al. (2024). Wu et al. (2020) showcased a U-Net Ronneberger et al. (2015)-based adversarial training framework which modifies input frames to decrease the accuracy of private attribute prediction while preserving action recognition performance. Dave et al. (2022b) proposed a self-supervised variant that focuses on reducing mutual information instead of relying on sensitive privacy labels.  Fioresi et al. (2023) adapts the self-supervised privacy objective from Dave et al. (2022b) for the anomaly detection task. Compared to the prior input-level anonymization methods our latent-space anonymization method differs in two key aspects: (1) previous methods are tailored to specific downstream tasks, such as action recognition in Dave et al. (2022b) and anomaly detection in Fioresi et al. (2023), while our approach aims to preserve privacy across various downstream video understanding tasks, (2) unlike prior methods, our method does not necessitate the retraining of the video model, thus providing computational efficiency for anonymizing even large-scale video foundation models.
@@
Bias Mitigation Computer vision tasks often struggle with spurious correlations Geirhos et al. (2018; 2020), where models rely on irrelevant information to make decisions, such as using background cues for action recognition instead of focusing on subjects’ movements Ding et al. (2022); Zou et al. (2023). Unfortunately, biases across a variety of protected demographic attributes, such as perceived gender, skin color, and age Zhao et al. (2017); Stock and Cisse (2017); Buolamwini and Gebru (2018); Wilson et al. (2019); Prabhu and Birhane (2020); Tong and Kagal (2020); Steed and Caliskan (2021); Gustafson et al. (2023); Narnaware et al. (2025); Kim et al. (2025); Swetha et al. (2025); Yousaf et al. (2025); Sirnam et al. (2026) have been found in vision-based tasks. These biases not only skew model performance but can also perpetuate harmful stereotypes. Barbano et al. (2021) explored the relationship between debiasing and privacy preservation, finding that there exists a subset of privacy preservation methods that are suitable for debiasing, giving promise to privacy preservation as a form of debiasing. In contrast to the image domain, biases in the video domain have not been as extensively studied. While a few papers Choi et al. (2019); Li et al. (2023a); Fioresi et al. (2025) address and mitigate scene bias in action recognition tasks, they overlook biases related to human attributes. Motivated by this gap, we introduce, for the first time, protocols to assess gender bias in action recognition. Our findings demonstrate that our self-supervised privacy preservation, even without an explicit bias-related objective, effectively generalizes in mitigating gender bias.
@@ 3.1 Problem Formulation
Privacy Preservation We propose a novel privacy-preserving framework that handles multiple utility tasks across diverse video datasets. Our framework is designed to maintain the high performance of a frozen video encoder across tasks while enforcing robust privacy constraints through the use of an embedding-level anonymization model $f_{A}$. Specifically, we consider video datasets that span action recognition ($\mathbb{D}_{AR}$), temporal action detection ($\mathbb{D}_{TAD}$), and anomaly detection ($\mathbb{D}_{AD}$). Each dataset $\mathbb{D}$ contains $N$ video samples, represented as $\{{\mathbf{x}^{(i)},\mathbf{y}^{(i)}\}_{i=1}^{N}}$, where $\mathbf{x}^{(i)}$ is a video instance, and $\mathbf{y}^{(i)}$ is its corresponding task-specific label. We define the set of utility tasks as $\{T_{AR};T_{TAD};T_{AD}\}\in T^{*}$. In implementation, we can choose a subset of tasks from $T^{*}$ for training, then evaluate on held-out tasks. We introduce a budget privacy evaluation task, denoted as $B$, where performance is measured by private attribute prediction. The framework starts with an off-the-shelf video encoder model $f_{E}$, left completely frozen. We define $f_{E}(\mathbf{x}_{t}^{(i)})=\mathbf{h}_{t}^{(i)}$ as the global clip-level embedding, taken as the [CLS] token for transformer encoders or the average-pooled feature for CNN-based encoders. In all experiments, $f_{A}$ operates only on this single clip-level embedding. The overall goal of $f_{A}$ is threefold: (1) to maintain the performance of $f_{E}$ across the set of defined utility tasks $T^{*}$, (2) to simultaneously reduce the performance on budget private attribute prediction task $B$, and (3) to preserve the general capabilities of $f_{E}$ such that performance is maintained on unseen tasks. This privacy preservation framework is outlined via the following criteria:
Criterion-1: Across each utility task, performance should be retained. Specifically, for task $T^{n}\in T^{*}$, the loss $\mathcal{L}_{T^{n}}$ before and after anonymization should be approximately equal.
$\begin{array}[]{l}\sum^{|T^{*}|}_{n}\mathcal{L}_{T^{n}}(f_{T^{n}}(f_{A}(f_{E}(X))),Y)\\ \approx\sum^{|T^{*}|}_{n}\mathcal{L}_{T^{n}}(f_{T^{n}}(f_{E}(X)),Y).\end{array}$ (1)
@@
Perceived Gender Bias Evaluation In the standard bias evaluation protocol, we are given a video dataset $\mathbb{D}_{AR}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}^{N_{IID}}_{i=1}$, where $\mathbf{x}^{(i)}$ is the $i$th video instance, $\mathbf{y}^{(i)}$ is the associated action label, and $N_{IID}$ is the number of dataset instances. After training, performance is evaluated on unseen bias test set $\mathbb{D}_{reco-OOD}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)})\}^{N_{OOD}}_{i=1}$, where $N_{OOD}$ is the number of out-of-distribution instances. The aim of any debiasing technique is to learn generalizable features of $\mathbb{D}_{IID}$ such that performance is maximized on $\mathbb{D}_{OOD}$ without compromising IID performance.
When considering perceived gender information, our in-distribution video dataset is now formulated as $\mathbb{D}_{AR}=\{(\mathbf{x}^{(i)},\mathbf{y}^{(i)},\mathbf{g}^{(i)})\}^{N_{IID}}_{i=1}$, where $\mathbf{g}^{(i)}\in\{male,female\}$ is the perceived binary gender label. We acknowledge that this binary formulation is not ideal and not inclusive of all gender categories. The bias evaluation test set also includes label $\mathbf{g}$ in order to evaluate subclass performance. Final evaluation is on a test set with a different label distribution.
@@ 3.2 Anonymization Framework
This section describes the full anonymization framework and training. The framework consists of 3 major components: (1) a frozen video encoder backbone $f_{E}$, (2) an anonymization function adapter $f_{A}$, which modifies the latent features while retaining the original shape, and (3) a set of utility classifier heads $\{f_{T_{AR}};f_{T_{TAD}};f_{T_{AD}}\}\in f_{T^{*}}$ for a predefined set of tasks. Note that $f_{A}$ acts solely on the global clip embedding $\mathbf{h}_{t}$ used as input to each utility task head.
Network Initialization To start, we initialize $f_{A}$ to act as an identity function. The video encoder model $f_{E}$ is initialized with off-the-shelf weights of Kinetics400 Carreira and Zisserman (2017) pretraining. Each $f_{T}$ classifier head matches a standard architecture for the provided task. For stability, these are initialized through non-anonymized training on their respective utility tasks. For action recognition, $f_{T_{AR}}$ is a simple linear layer. For action detection $f_{T_{TAD}}$ and anomaly detection $f_{T_{AD}}$ architectures, TriDet Shi et al. (2023) and MGFN Chen et al. (2023) respectively are utilized.
Anonymization Training The training process consists of an adversarial optimization between a budget privacy loss $\mathcal{L}_{B}$ and a collection of standard utility losses $\mathcal{L}_{T^{*}}$, regularized by a proposed latent consistency loss $\mathcal{L}_{LC}$.
@@
Clip-Level Budget Privacy Objective Our clip-level self-supervised budget privacy objective is the key component for facilitating anonymization without requiring private attribute labels. The intuition is that two frames share a lot of mutual information, so if we minimize the similarity between them, the shared spatial information gets destroyed. A crucial difference setting SPLAVU apart from prior works is that the anonymizer works across the temporal dimension using 3D clip features instead of a 2D U-Net Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023). This way, when combined with utility task losses, the anonymization model learns to remove all spatial information, maintaining only temporal information useful for solving the utility task. We utilize the SimCLR NT-Xent Chen et al. (2020) contrastive loss as $\mathcal{L}_{B}$, defined as follows:
@@
where $\mathbf{\bar{h}}_{t}^{(i)}$ represents the feature vector of a static clip sampled from video $\mathbf{x}^{(i)}$ at time $t$, $d(u,v)=exp(u^{T}v/(\|u\|\|v\|\tau))$ computes the similarity between the input vectors with temperature parameter $\tau$. $\mathds{1}_{[j\neq i]}$ is an indicator function that equals 1 when $j\neq i$. Minimizing this loss increases the similarity between inputs $\mathbf{\bar{h}}_{\bar{t_{1}}}^{(i)}$ and $\mathbf{\bar{h}}_{\bar{t_{2}}}^{(i)}$ over the sum of all other clips in the batch. Instead, we opt to maximize the loss, resulting in the objective destroying mutual information between these clips instead. Notably, as opposed to previous works Wu et al. (2020); Dave et al. (2022b); Fioresi et al. (2023), we do not use a disjoint image encoder model to compute $\mathbf{\bar{h}}_{\bar{t}}^{(i)}$. Alternatively, the video encoder model $f_{E}$ itself is used to process clip frames (tiled to match a standard clip shape, see Figure 2), leading to a much more natural utility-privacy interaction for improved anonymization. These static clip features are then directly utilized in this budget privacy loss (Eq. 5).
@@
where $\omega_{LC}$, $\omega_{T}$, and $\omega_{B}$ are weights to control the strength of each objective. The privacy term is maximized via the negative sign in $\omega_{B}*\mathcal{L}_{B}$ and works against $\mathcal{L}_{T}$ and $\mathcal{L}_{LC}$ in a GAN-style paradigm, until the anonymizer is able to remove all encoded spatial information except for what is necessary for performance on the utility tasks. After this training, we are left with a lightweight anonymization adapter $f_{A}$ that can be appended to the off-the-shelf video encoder model $f_{E}$ for use in a variety of downstream tasks.